OpenStudy (anonymous):

what is the answer to w+ 5/6=7

OpenStudy (anonymous):

w + 5 over 6 = 7

OpenStudy (anonymous):

what is the answer to w+5\6 =7

OpenStudy (anonymous):

no one knows the answer

OpenStudy (anonymous):

\[w+\frac{ 5 }{ 6 } = 7\] \[w = 7 - \frac{ 5 }{ 6 }\] \[w = \frac{ 37 }{ 6 }\]